# Benchmarking Multi-Agent Preference-based Reinforcement Learning for Human-AI Teaming

### v1

@@ 6 Benchmarking Human-AI Team PbRL
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]
@@
6.1 Evaluation Domains
6.1.1 Multi-Agent Highway Orchestration Task:
We first design a set of three multi-agent environments built on top of the Highway environment [11]. As shown in Figure 2, we construct a three lane highway where the human car (in orange) is on the top-most lane, and the AI agent (in purple) has to overtake and merge back into the human car’s lane by increasing its speed and avoiding collisions with any other car in the scene. The human prefers that the AI agent car overtakes from the right-most (bottom-most in the figure) lane and not from the middle lane. Similarly, we construct two more scenarios. In the second one, both the cars are in the right lane and the human preference is for the AI agent car to overtake from the left-most lane. In the third domain, both cars are in the middle lane and the AI agent can only overtake from the left lane. In the results shown in Figure 4, these environments are referred to as MA Highway - Right Overtake, MA Highway - Left Overtake, and MA Highway - Middle Overtake respectively.
6.1.2 Multi-Agent Locomotion Orchestration Tasks:
We present a more complex domain suite, which necessitates cooperation, to study Human-AI team PbRL cooperation game. Following prior works in constructing multi-agent environments [55] in continuous control domains, we modify the MuJoCo-based locomotion environments [56] for the human-AI teaming tasks. The locomotion tasks include a robot attempting to move (for example by a walk, run or crawl). Such tasks are essentially tacit knowledge robotic simulation tasks for which reward specification can be challenging [14, 9]. Moreover, single-agent variant of these environments have been used as benchmark domains by single-agent PbRL works [14, 13, 15], further motivating their use.
We construct multi-agent settings for Cheetah, Ant, Walker, Swimmer & Hopper environments, as shown in Figure 3. In each environment, the human agent controls the orange highlighted joints action space, while the rest of the joint angle controls are for the AI agent highlighted in purple. For example, in the MA-Walker, the human controls the walker’s right leg joints to walk forward, while the AI agent controls the left leg joints.
Figure 4: Learning curves on MA Highway (rows 1 to 3) and MA MuJoCo - Cheetah (row 4): comparing (from L to R) (a) 0% access to $\pi_{H}$, (b) partial access to $\pi_{H}$, and (c) human flexibility on multiple $\pi_{H}$, with Specified Orchestration case that assumes a single $\pi_{H}$ and complete access to it; as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
((a))
((b))
((c))
((d))
Figure 5: Learning curves on MA MuJoCo domains: comparing PbRL algorithms extended to the multi-agent setting under Specified Orchestration, as measured on the episodic returns. The solid lines and shaded regions represent the mean and standard deviation, respectively, across three runs.
6.2 Adapting Single Agent PbRL to Human-AI Team setting
We present our modified PbRL algorithm for the human-AI team setting, which uses the PEBBLE algorithm [15] as a backbone. Note, we make the same modifications for experiments using RUNE [13] and SURF [14] as these algorithms also use PEBBLE as their PbRL backbone. We do not require any change to the model architectures (including reward approximator and policy approximator). The updates are highlighted in red (see Algorithm 1) and can be easily incorporated into any online single agent PbRL algorithm.
Specifically, we assume a human policy set $|\Pi|$ from which the human agent can select a single (or multiple) policies for themselves, and the AI agent has the access to the human policy during training with a budget $b$. When a new episode begins, we uniformly sample a human policy $\pi_{H}$ from $|\Pi|$. Note, if the human only has a single policy, i.e., in the case of Specified Orchestration, then we use the same $\pi_{H}$ for every episode. During the policy learning phase for the AI agent if the agent has remaining budget to request actions from human policy, we sample $a_{H}$ from $\pi_{H}$, or else, the AI agent assumes a random action on behalf of the human. Again, in the case of Specified Orchestration, the budget $b$ allows the AI agent to access human actions till the end of the training. As per our re-interpretation of the Human-AI team PbRL Cooperation game in Section 5.1, this is equivalent to the assumption that human is acting in the shared world with the agent.
Following past works in Preference based Reinforcement Learning, we use a simulated oracular human policy to provide the actions on behalf of the human agent. As suggested in [22] we use a scripted preference labeling function that provides preference labels on behalf of the human agent to the PbRL agent. We train SAC [57] by assuming centralized controller for human and AI agent using underlying task reward available with the environment package for MuJoCo [12] and handcrafted reward function for Highway. We take the actions corresponding to the joint angles under control of the human as human actions and rest for the AI agent’s actions expected by the human as part of the compatible policy pair. Preference labels are generated based on hand-crafted rewards for MA Highway and environment packaged rewards (which serve as team reward) for MA MuJoCo. We refer the readers to the appendix for further details.
6.3 Impact of Access to Human Policy
6.3.1 Specified Orchestration - complete access to $\pi_{H}$:
To answer Question (2) and to support our discussion on Specified Orchestration, we report the performance of all three PbRL algorithms on the three MA Highway, and five MA MuJoCo environments in Figures 4 and 5. We note from each of the plots, that PEBBLE [15], RUNE [13] and SURF [14] show the optimal returns. From Figure 5, we also note that RUNE consistently performs better than SURF and PEBBLE in these multi-agent settings.
6.3.2 Restricting Access to Human Policy:
To answer Question (3), we first investigate the case where we restrict the AI agent’s budget to access the human policy $\pi_{H}$ to 50% for MA Highway domains, and 10% for MA MuJoCo - Cheetah. We allocate this entire budget in the first half of the training. For example, in the case of MA MuJoCo - Cheetah domain, the AI agent has access to $\pi_{H}$ up-till 100k steps, and assumes random actions sampled for the remaining training steps. We report the results in the second column of Figure 4. Surprisingly, we note that the performance plummets significantly as soon as the human agent’s actions are not available for all the domains.
6.3.3 No Access to Human policy:
In reference to Question (3), we report the performance of PEBBLE [15], RUNE [13], and SURF [14] on the three MA Highway and MA MuJoCo - Cheetah domains in first column of Figure 4. None of the algorithms have any access to $\pi_{H}$, and thus, have to ‘imagine’, i.e., assume random actions for the human agent. In all the four domains and three algorithms, we note that the performance is significantly worse as compared to the case when the AI agent has complete access to $\pi_{H}$. Utilizing explicit modeling of human policy is a natural next step and an important future work, but it beyond the scope of the first exploration of PbRL in a Multi-agent setting.
6.3.4 Observed Challenges:
From the experiments with 0% and partially restricted access to the human partner policy, we observe that reward learning is severely impacted. Since Specified Orchestration assumes maximal information on the human partner’s policy and preferences, the team performance is optimal. This highlights the potential of using Specified Orchestration as a means to obtain an loose upper bound on the Human-AI team PbRL cooperation game performance. Further, unless the AI agent explicitly models human policy (and preferences) with a restricted access, reward learning along with accomplishing the team task stays a challenging problem.
6.4 Impact of Human Flexibility
6.4.1 Multiple Human policies:
We observe that each of three PbRL approaches perform the best in the case where the agent had complete access to $\pi_{H}$. To study the impact of Human Flexibility (our Q1) on PbRL performance we operate under the assumption that agent has complete access to the human policy (i.e. human executes their policy during agent’s training)
We assume that for each episode of learning, the human agent can sample 1-out-of-N (3, in our case) policies to provide the actions to the AI agent. From the results shown in the third column of Figure 4, we note that the performance comparatively drops as in the case where we assumed a single $\pi_{H}$ and complete access to it. The decline in performance is amplified in the more complex case of MA MuJoCo Cheetah where cooperation is forced amongst the two agents. We conclude that the team performance is the highest in the case of Specified Orchestration, where the human agent has a single $\pi_{H}$ and the AI agent has complete access to the same.
6.4.2 Observed Challenges:
7 Conclusion & Future Work
We perform a first exploration of Preference-based Reinforcement Learning for Human-AI team setup. We introduce a Human-AI PbRL Cooperation game as a special case of Cooperative Inverse Reinforcement Learning [10] and argue across two critical dimensions, namely - AI agent’s Access to Human Policy and Human Flexibility. In our heterogeneous Human-AI cooperative setup, both the agents are rewarded according to human’s reward function for the team comprising of human’s objectives for the team, sub-objectives for themselves and the robot, and human preferences on team behavior. For investigating the performance of PbRL approaches, we modify the existing state-of-the-art single-agent PbRL algorithms (namely, PEBBLE [15], SURF [14] & RUNE [13]). To this end, we benchmark PbRL algorithms on our benchmark suite consisting of simpler Highway domains and more complex Robot locomotion.
We note that all three PbRL algorithms gave optimal returns only when the human partner has a preference for a single policy pair (team behavior) and that they have complete access to human partner’s policy (termed as Specified Orchestration). We discuss how Specified Orchestration can be useful to benchmark future works in Human-AI team PbRL as a loose upper bound on algorithm performance. To our surprise, partial access to $\pi_{H}$ also had a drop in performance as soon as the agent’s access budget was exhausted. Subsequently, we also observe poor performance when the AI agent does not have access to human partner’s policy. We conclude that there is a need for AI agents to explicitly model partner human policy (along with their preferences). Furthermore, utilizing these observations, we assumed complete access to $\pi_{H}$, and ablated in the case where the human partner is flexible, i.e., has multiple acceptable team strategies. Counter-intuitively, we find that when the human is more flexible, the agent’s performance declines. This is primarily because we constrain our AI agent to learn a conformant policy (compatible with all human policies).
From our benchmarking results, we conclude that Specified Orchestration, while assuming the maximal amount of information on the human, shows notable performance across all domains. However, this can also be seen as one of the key limitations as it may not always be the case that the human can easily compute and provide their policy to the AI agent (or be available to act during the training procedure of the AI agent). Secondly, while this approach allows for optimal team performance of an AI agent with a specific human, it may be hard to generalize the learnt reward function (and hence, the learnt policy) to multiple human partners. Lastly, as in the case of single-agent PbRL works where human is just an observer, the problem of expensive human interaction with the AI agent is carried over to our setting. To conclude, our work provides researchers with a first exploration of PbRL beyond single-agent setup, towards human-AI teaming.
8 Societal Impact
Appendix A PbRL Algorithm
A.1 PEBBLE Algorithm
@@
PEBBLE is a PbRL algorithm that comprises of two key elements: pre-training and relabeling experience buffer. To gather a wide range of experiences, PEBBLE starts by using intrinsic motivation [58, 59, 60, 61] to pre-train the policy, which optimizes the policy to increase the state entropy in order to explore the environment effectively. Afterwards, PEBBLE uses the SAC algorithm, a state-of-the-art off-policy RL algorithm, to further train the policy. To ensure stability in the learning process, PEBBLE relabels all experiences in the buffer when the reward model is updated.
@@
A.2 Modified PEBBLE Algorithm for varying Access to Human Policy experiments
@@
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]
@@
A.3 Modified PEBBLE Algorithm for Human Flexibility experiments
@@
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in [15]
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in [15]
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively [15]

### v2

@@ Appendix A PbRL Algorithm
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in (Lee, Smith, and Abbeel 2021)
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in (Lee, Smith, and Abbeel 2021)
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively (Lee, Smith, and Abbeel 2021)
@@
PEBBLE Algorithm
@@
PEBBLE is a PbRL algorithm that comprises of two key elements: pre-training and relabeling experience buffer. To gather a wide range of experiences, PEBBLE starts by using intrinsic motivation (Chentanez, Barto, and Singh 2004; Barto 2013; Abel et al. 2021; Schmidhuber 2010) to pre-train the policy, which optimizes the policy to increase the state entropy in order to explore the environment effectively. Afterwards, PEBBLE uses the SAC algorithm, a state-of-the-art off-policy RL algorithm, to further train the policy. To ensure stability in the learning process, PEBBLE relabels all experiences in the buffer when the reward model is updated.
@@
Modified PEBBLE Algorithm for varying Access to Human Policy experiments
Please refer to Algorithm 2 for more details.
@@
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in (Lee, Smith, and Abbeel 2021)
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in (Lee, Smith, and Abbeel 2021)
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively (Lee, Smith, and Abbeel 2021)
@@
Modified PEBBLE Algorithm for Human Flexibility experiments
Please refer to Algorithm 3 for more details.
@@
  $D_{\tau},\pi_{\phi}\leftarrow$ EXPLORE() in (Lee, Smith, and Abbeel 2021)
@@
           $(\sigma^{0},\sigma^{1})\sim$ SAMPLE() in (Lee, Smith, and Abbeel 2021)
@@
        Store transitions $D_{\tau}\leftarrow D_{\tau}\{(s_{t},a_{t},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}a_{h}},s_{t+1},\hat{r}_{\psi}(s_{t}))\}$
@@
        Optimize $L^{SAC}_{critic}$ and $L^{SAC}_{actor}$ w.r.t. $\theta$ and $\phi$, respectively (Lee, Smith, and Abbeel 2021)
@@
Appendix B User Study Details
All participants were provided with a detailed information page explaining the study’s instructions, expected duration and payment details. Participants were assured that their involvement in the study adheres to strict confidentiality measures, aligning with IRB guidelines (code redacted for anonymity). By continuing with the study, participants consented to the secure recording of their data for research purposes.
Figure 6: Welcome Page for the Use Study
Figure 7: Description of the Task Objectives, Controls and Legend
Figure 8: Example Videos to Illustrate Multiple Joint Policies for Task Success and practice game with controls.
Figure 9: Instructions Provided for Adaptive Case
Figure 10: Instructions Provided for Specified Orchestration Case
Study Design
Inspired by the Overcooked domain (Carroll et al. 2019), we implement a multi-player cooperative game where participants engage with a Helper Agent in an environment featuring Onion dispensers, Dish dispensers, Oven, a Serving station, and empty counters, with the goal of cooking and delivering onion soup dishes. The dynamic game-play allows players to move within the environment, executing actions such as picking, placing, and collecting objects. The process of preparing onion soup involves picking onions, placing them in the oven, and collecting the soup in a dish. Successful delivery of the soup to the pickup counter rewards both players with 20 points. The study introduces different treatments within the same control group, varying the sizes of policy sets to observe the impact on subjective experience and team performance. Participants initially come across a page containing game instructions. This page provides details on game controls, task objectives, and a legend explaining the symbols and their corresponding meanings in the game environment as shown in Figure 7. Next, participants are presented with demo videos illustrating examples of how the task can be accomplished by the two agents with different levels of collaboration and allowed to test out the controls in the game environment (Figure 8). In the first treatment, participants have the flexibility to select from a large set of policies, adapting to the Helper Agent’s dynamic execution of actions during the game (Figure 9). The instructions provided to the agents in the webpage are - ”The helper agent is executing certain actions and you are expected to adapt and perform actions accordingly. For example, in the game, Suppose the helper agent moves left toward the pick-up counter, You need to deliver onion soup to the pick-up counter. The actions available to you at each step are provided in the interface below the grid. You can choose any action available to you”. The second treatment introduces a fixed policy for the Helper Agent, wherein players receive a predetermined plan to follow (Figure 10). This is the case of Specified Orchestration where the set of feasible strategies is singleton for the players. Players are required to adhere to their chosen policy at each step, helping investigate how different policy set sizes influence the overall game-play experience and team dynamics. The following instructions are provided to the participants - ”The Helper Agent has a fixed policy (The Agent has a fixed plan and will execute actions in a fixed predetermined order). You are given a plan that you can follow. And you are required to execute actions according to that plan. Please stick to your chosen policy when executing your actions at each step. The suggested action relating to your chosen policy will be highlighted in red color”.
Additional Results
Demographics Statistics:
The user study involves a diverse demographic, with participants averaging 35 years of age and exhibiting a standard deviation of 12.78. The gender distribution showcases 6 male, 13 female, and 1 participant identifying as ’Other,’ ensuring a varied and representative sample for the study.
Table 1: User Study Demographics Statistics: The User study respondents consisted of 20 Participants with the following age and gender distributions.
Age Statistics
Average Standard Deviation
$35$ $12.78$
Gender Statistics
Male Female Other
$6$ $13$ $1$
Session Time:
In the adaptive case, the total session time was 2 minutes and 1 second, surpassing 1 minute and 16 seconds recorded in the rigid (Specified Orchestration) case. Interestingly, even though users in the rigid case invested time upfront in policy selection before the run begun, the total active session time (the duration when users actively engaged in the task) was markedly shorter than that observed in the adaptive case. This reinforces the argument for Specified Orchestration, where the human communicates it’s selected policy to the agent, and maintains a single policy in their set of feasible team strategies.
We also conducted the Student’s t-test to compare the total session time for all participants in the rigid vs. adaptive case. Our null hypothesis is: there is no difference between the time taken by participants to finish the task in the rigid (Specified Orchestration) case and the adaptive case. The t-test revealed that there is a statistically significant difference in the total session time (p = 0.043073 $<$ 0.05) as well as the total active session time (p = 0.02953 $<$ 0.05) for participants in the rigid case and adaptive case. Thus, we reject the null hypothesis.
Table 2: User Study Time Metrics
Time Metrics Case Time (s)
Average Total Session Time Specified Orchestration 81
Adaptive Case 112
Average Active Session Time Specified Orchestration 52
Adaptive Case 82
Subjective Metrics:
At the end of the user study, we administer a questionnaire to assess subjective aspects of the user experience for the game, specifically focusing on mental demand, time pressure, task efficiency, and frustration for both cases. These factors directly contribute to the evaluation of cognitive load, temporal load, task performance and frustration as shown in Table 3.
Questionnaire:
  1. 1.
    Relative Mental Demand:
    1. (a)
      How mentally demanding was completing the overall task, relatively in the two cases?
    2. (b)
      How mentally demanding was generating your actions, relatively in the two cases?
    3. (c)
      How mentally demanding was predicting the actions of the other agent at every step, relatively in the two cases?
  2. 2.
    Relative Time Pressure: How hurried or rushed did you feel to come up with your next action at every step, relatively in the two cases?
  3. 3.
    Relative Efficiency:
    1. (a)
      How efficient do you think you were in accomplishing the task, relatively in the two cases?
    2. (b)
      How efficient do you think you were in cooperating with the other agent to complete the task, relatively in the two cases? (Cooperating here refers to the ease of working together without any direct communication with your partner.)
    3. (c)
      How efficient do you think you were in predicting what the other agent is going to do at every step, relatively in the two cases?
  4. 4.
    Relative Frustration: How insecure, discouraged, irritated, stressed, and annoyed were you, relatively in the two cases?
  5. 5.
    Frustration for Adaptive Case: On a scale of 1-5, rate your frustration for the following four parameters in Case 1:
    1. (a)
      Deciding your next action:
    2. (b)
      Understanding the other agent’s next action or intention:
    3. (c)
      Adapting to the other agent to ensure task completion:
    4. (d)
      Relying on the other agent to ensure task completion:
  6. 6.
    Frustration for Rigid Case: On a scale of 1-5, rate your frustration for the following four parameters in Case 2:
    1. (a)
      Deciding your next action:
    2. (b)
      Understanding the other agent’s next action or intention:
    3. (c)
      Adapting to the other agent to ensure task completion:
    4. (d)
      Relying on the other agent to ensure task completion:
To gauge relative user preferences between the adaptive and rigid (Specified Orchestration) scenarios, participants provide input on a Likert scale ranging from 1 (’Most in Case 1’) to 7 (’Most in Case 2’), with 4 indicating equality in both cases. Additionally, frustration levels are evaluated on a separate Likert scale, ranging from 1 (’Minimum frustration’) to 5 (’Maximum frustration’), aiming to comprehensively understand and compare the frustration experience by the users in both cases. Statistics for participants’ responses to the questions are presented in the Figure 3 and visualized in Figure 11, Figure 12, Figure 13, Figure 14, Figure 15, Figure 16, Figure 17 and Figure 18 to offer a comprehensive overview of the subjective aspects, including mental demand, time pressure, task efficiency, and frustration. The provided statistical plots offer a comprehensive view of the participants’ experiences across various dimensions including the metrics from Table 3. Participants reported a significantly higher mental demand in the adaptive case as compared to the specified orchestration case for completing overall task, generating their actions and predicting other agents’ actions. The urgency to generate the next action indicated a notable time pressure for the adaptive case, suggesting a pressure to generate actions in real-time for users in this case. Regarding task efficiency, respondents perceived higher efficiency in completing tasks and cooperation in the Specified Orchestration (rigid) case, where predictive abilities regarding the other agent’s actions were notably high. Frustration levels were higher in the adaptive case, with the adaptive case showing higher frustration in decision-making, understanding the other agent’s intention, and reliance on the other agent. Conversely, specified orchestration demonstrated lower frustration levels across these parameters.
Table 3: User Study Descriptive Statistics for Evaluation
Evaluation Factors Study sub-factors Question Focus Mean Standard Deviation Median
Cognitive Load Relative Mental Demand Completing the overall task 3.1 1.74 3
Generating your actions 3.05 1.85 3
Predicting Other Agent’s actions 3.3 1.72 3
Temporal Load Relative Time Pressure Urgency to generate next action 3.25 1.55 4
Task Performance Relative Task Efficiency Efficiency in completing Task 3.95 2.09 4
Efficiency in cooperation 4.4 1.73 4.5
Predicting Other Agent’s actions 4.65 1.6 4
Frustration Relative Frustration/ Stress Relative Frustration or Stress 3.35 1.73 3
Frustration - Adaptive Case Deciding next action 2.75 1.59 2
Understanding other agent’s Intention 2.95 1.57 2
Adapting to other agent 3 1.34 3.5
Relying on the other agent 3 1.52 3
Frustration - Specified Orchestration Deciding next action 2.2 1.4 1.5
Understanding other agent’s Intention 2.6 1.35 2.5
Adapting to other agent 2.7 1.49 3
Relying on the other agent 2.6 1.6 2
Figure 11: Boxplots representing Cognitive Load factors that shows responses from the User Study evaluating relative mental demand on a Likert Scale of 1 to 7 for (i) Completing the Overall task. (ii) Generating their actions. (iii) Predicting Other Agent’s actions. The green and orange lines within the boxes represent the mean and median respectively.
Figure 12: Boxplots representing Temporal Load responses from the User Study evaluating relative task pressure on a Likert Scale of 1 to 7. The green line within the box represents the mean.
Figure 13: Boxplots representing Task Performance factors that show subjective responses from the User Study evaluating relative task efficiency on a Likert Scale of 1 to 7 for (i) Efficiency in Task Completion. (ii) Efficiency in Cooperation. (iii) Predicting Other Agent’s actions. The green and orange lines within the boxes represent the mean and median, respectively.
Figure 14: Boxplots representing Relative Frustration and Stress that show subjective responses from the User Study on a Likert Scale of 1 to 7. The green and orange lines within the box represent the mean and median, respectively.
Figure 15: Boxplots comparing User Frustration levels when deciding their next action in Specified Orchestration and Adaptive Cases using a Likert Scale of 1 to 5 as shown. The green and orange lines within the boxes represent the mean and median, respectively.
Figure 16: Boxplots comparing User Frustration levels in understanding other agent’s intention for Specified Orchestration and Adaptive Cases using a Likert Scale of 1 to 5 as shown. The green and orange lines within the boxes represent the mean and median, respectively.
Figure 17: Boxplots comparing User Frustration levels when adapting to the other agent in Specified Orchestration and Adaptive Cases using a Likert Scale of 1 to 5 as shown. The green and orange lines within the boxes represent the mean and median, respectively.
Figure 18: Boxplots comparing User Frustration levels when relying on the other agent to ensure task completion in Specified Orchestration and Adaptive Cases using a Likert Scale of 1 to 5 as shown. The green and orange lines within the boxes represent the mean and median respectively.
Appendix C Experiment Details
Evaluation Domains
Multi-Agent Highway Orchestration Task:
We first design a set of three multi-agent environments built on top of the Highway environment (Leurent 2018). As shown in Figure 2, we construct a three lane highway where the human car (in orange) is on the top-most lane, and the AI agent (in purple) has to overtake and merge back into the human car’s lane by increasing its speed and avoiding collisions with any other car in the scene. The human prefers that the AI agent car overtakes from the right-most (bottom-most in the figure) lane and not from the middle lane. Similarly, we construct two more scenarios with both cars in the right and middle lanes. In Figure 4, these are referred to as MA Highway - Right Overtake, MA Highway - Left Overtake, and MA Highway - Middle Overtake, respectively.
Multi-Agent Locomotion Orchestration Tasks:
Following prior works in constructing multi-agent environments (Peng et al. 2021) in continuous control domains, we modify the MuJoCo-based locomotion environments (Todorov, Erez, and Tassa 2012) for the human-AI teaming tasks which necessitate cooperation. Single-agent variant of these environments have been used as benchmark domains by single-agent PbRL works (Park et al. 2022; Liang et al. 2022; Lee, Smith, and Abbeel 2021), further motivating their use. We construct multi-agent settings for Cheetah, Ant, Walker, Swimmer & Hopper environments, as shown in Figure 3. In each environment, the human agent controls the orange highlighted joints action space, while the rest of the joint angle controls are for the AI agent highlighted in purple. For example, in the MA-Walker, the human controls the walker’s right leg joints to walk forward, while the AI agent controls the left leg joints.
State Space and Action Space